Solve the differential equations y"-3y'+2y=0
1
Expert's answer
2019-09-10T10:43:04-0400

It's a second-order linear ordinary differential equation.

An equation "y''+py'+qy=0" has solution "y=C_1e^{\lambda_1x}+C_2e^{\lambda_3x}" where "\lambda_1,\lambda_2" are the roots of "x^2+px+q=0" if "\lambda_1\ne\lambda_2" and "y=(C_1+C_2x)e^\lambda" if "\lambda_1=\lambda_2\ (=\lambda)" .

https://en.wikipedia.org/wiki/Linear_differential_equation#Second-order_case

So, "x^2-3x+2=0" has the roots "1" and "2" , according to converse Vieta's theorem (as "-(1+2)=-3" and "1\cdot 2=2" ).

So, the solution is "y=C_1e^x+C_2e^{2x}" .

You can see here https://www.wolframalpha.com/input/?i=y%22-3y%27%2B2y%3D0 , it's correct.